DCCC M-121 (SP18) Quiz 8 04/25/2018 1) The amount of zinc in vitamins (in PPM) was determined for 27 samples, resulting in the following data. The data values have been arranged in order from largest to smallest 30 32 6 63 70 79 87 90 101 136 125 279 120 119 198 118 115 182 118 183 118 190 102 195 172 a) Create a stem and leaf plot (use the stems below): 10 12 13 14 15 16 17 18 19 20 21 23 24 25 26 27 28 29 31 32 34 35
media%2Fc75%2Fc750e494-6b32-4e77-9db6-44
Answer #1
  • stem leaves
    3 0,2
    6 0,3
    7 0,9
    8 7
    9 0
    10 1,2
    11 5,8,8,8,9
    12 0.5
    13 6
    17 2
    18 2,3,
    19 0,5,8
    20 0
    27 9
    35 9
  • The formula for the mean is given by \bar{X}=\sum X/n =134.85
  • The median of the data set id given by n/2+1/2 the value 27/2+1/2=14th value. i.e 118
  • The mode is the value which occurs most often i.e 118
  • The minimum value is 30 and max value is 359.
  • The Range of the value is Maximum -Minimum=259-30=329
  • May 35 360 33. 2 2.9 o 2.10 edian c ower 60 3
  • AS above we can see the boxplot the middle line is median which is 118 the upper quartile is 182.5 and lower quartile is 88.5 .